# Robustness of topological phases on aperiodic lattices

### v1

@@ 2. Topological phases and their numerical invariants
We recall the definition of “real” and real C*-algebras. In other literature, a “real” C*-algebra is also referred to as a Real C*-algebra (cf. [Bourne-Kellendonk-Rennie:Cayley_transform], note the upper case R) or a $\mathrm{C}^{*,\mathrm{r}}$-algebra (cf. [Kellendonk:Cst-algebraic_topological_phases]).
@@
When there is no ambiguity, we shall use $A$ alone to denote the “real” C*-algebra $(A,\mathfrak{r})$. For example, if $A=\mathbb{M}_{n}\mathbb{C}$ is the C*-algebra of complex $n\times n$-matrices, then the entrywise complex conjugation defines a “real” structure on $A$.
Definition 2.2.
A “real” involution on a $\mathbb{Z}/2$-graded Hilbert space $\mathcal{H}$ is a conjugate-linear, grading-preserving automorphism of $\mathcal{H}$ of order two.
Let $\Theta$ be a “real” involution on a $\mathbb{Z}/2$-graded Hilbert space. Then $\mathbb{B}(\mathcal{H})$ becomes a $\mathbb{Z}/2$-graded “real” C*-algebra with “real” structure $T\mapsto\Theta T\Theta$.
Definition 2.3.
@@
Now we define real C*-algebras and real Hilbert spaces.
@@
If $A$ is a “real” C*-algebra, then the collection of its real elements
@@
2.2. Clifford algebras
Let $\mathbb{C}\ell_{p,q}$ be the finite-dimensional graded “real” C*-algebra generated by $\gamma^{1},\dots,\gamma^{p},\rho^{1},\dots,\rho^{q}$, satisfying:
@@ 2.2. Clifford algebras
Next, we describe the representation of the graded “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ following [Ewert-Meyer:Coarse_geometry]*Section 4. The grading on $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ is the tensor product of the trivial grading on $\mathrm{C}^{*}(\mathbb{Z}^{d})$ with the standard grading on $\mathbb{C}\ell_{0,d}$. The “real” structure on $\mathrm{C}^{*}(\mathbb{Z}^{d})$ is given by the pointwise complex conjugation, that is,
$\mathfrak{r}(f)(n)\mathrel{\vcentcolon=}\overline{f(n)},\quad n\in\mathbb{Z}^{d}.$
@@
The Fourier transform maps the “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})$ to the “real” C*-algebra $\mathrm{C}(\mathbb{T}^{d})$. Here $\mathbb{T}^{d}$ is a “real” manifold, with involution $z\mapsto\overline{z}$.
@@
$\left\lvert k\right\rangle\otimes e_{i_{1}}\wedge\dots e_{i_{l}}\mapsto\frac{z^{k}}{z_{i_{1}}\dots z_{i_{l}}}\mathrm{d}z_{i_{1}}\wedge\dots\wedge\mathrm{d}z_{i_{l}},$
@@
$\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}\simeq\mathrm{C}(\mathbb{T}^{d})\otimes\mathbb{C}\ell_{0,d}\simeq\mathrm{C}(\mathbb{T},\mathbb{C}\ell(\mathbb{T}^{d}))$
where $\mathbb{C}\ell(\mathbb{T}^{d})$ is the “real” Clifford bundle of the “real” manifold $\mathbb{T}^{d}$. In particular, under this isomorphism, the canonical representation of $\mathbb{C}\ell_{0,d}$ as in 2.5 is mapped to the Clifford multiplication of $\mathbb{C}\ell(\mathbb{T}^{d})$ on $L^{2}\left(\bigwedge\nolimits^{*}\mathrm{T}^{*}\mathbb{T}^{d}\right)$.
Let ${\scriptstyle\mathsf{X}}_{j}$ be the $j$-th self-adjoint position operator on $\ell^{2}(\mathbb{Z}^{d})$ defined by
${\scriptstyle\mathsf{X}}_{j}f(k)\mathrel{\vcentcolon=}k_{j}f(k)$
with domain the space of rapidly decaying functions $\mathbb{Z}^{d}\to\mathbb{C}$. Thus the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ is odd, self-adjoint acting on the $\mathbb{Z}/2$-graded “real” Hilbert space $\ell^{2}(\mathbb{Z}^{d})\otimes\bigwedge\nolimits^{*}\mathbb{C}^{d}$.
@@
by the standard representation of $\mathbb{C}\ell_{d,0}$ in 2.5. The generators $\gamma^{j}$ are “real” in $\mathbb{C}\ell_{d,0}$, hence maps $\bigwedge\nolimits^{*}\mathbb{R}^{d}$ to $\bigwedge\nolimits^{*}\mathbb{R}^{d}$. The $j$-th position operator ${\scriptstyle\mathsf{X}}_{j}$ preserves $\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}$ because ${\scriptstyle\mathsf{X}}_{j}\left\lvert k\right\rangle=k_{j}\left\lvert k\right\rangle$ and $k_{j}$ is real for any $k\in\mathbb{Z}^{d}$. Therefore the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ preserves $\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$. ∎
2.3. Topological phases and K-theory
We briefly recall the noncommutative framework for symmetry-protected topological phases of matter. Let $\mathcal{H}$ be a complex Hilbert space, which we regard as the “physically relevant” one. The $\mathsf{CT}$-symmetry of a physical system is described by:
  • a finite group $G\subseteq\mathbb{Z}/2\times\mathbb{Z}/2$;
  • a group homomorphism $\varphi\colon G\to\{\pm 1\}$;
  • a 2-cocycle $\sigma\colon G\times G\to\mathbb{T}$;
  • a Borel map $g\mapsto\theta_{g}$ from $G$ to the set of maps on $\mathcal{H}$
such that
  • $\theta_{g}$ is a unitary (resp. anti-unitary) operator on $\mathcal{H}$ if $\varphi(g)=+1$ (resp.$\varphi(g)=-1$ );
  • $\theta_{g_{1}}\theta_{g_{2}}=\sigma(g_{1},g_{2})\theta_{g_{1}g_{2}}$.
The tuple $(G,\varphi,\sigma,\theta)$ is called a projective unitary/anti-unitary (PUA) representation of the $\mathsf{CT}$-symmetry group $G$ , cf. [Thiang:K-theoretic_classification, Bourne-Carey-Rennie:NC_framework_topological_insulators].
Let $A$ be a $\mathbb{Z}/2$-graded “real” C*-subalgebra of $\mathbb{B}(\mathcal{H})$ An abstract insulator in $A$ is a self-adjoint element $h\in A$ which has a spectral gap at $\mu$, and such that $h$ is compatible with the $\mathsf{CT}$-symmetry group. The characteristic function of $(-\infty,\mu)$ is continuous on the spectrum of $h$, thus continuous functional calculus gives an element $p_{\mu}\mathrel{\vcentcolon=}\chi_{(-\infty,\mu)}(h)$, called the Fermi projection. The following proposition comes from [Kellendonk:Cst-algebraic_topological_phases]*Section 6.
Proposition 2.9.
Let $h$ be a self-adjoint element in a $\mathbb{Z}/2$-graded “real” C*-algebra $(A,\mathfrak{r})$, which has a spectral gap at $\mu$. If $h$ has no symmetry, then $h-\mu$ represents a class in $\operatorname{K}_{0}(A)$, given by its Fermi projection $p_{\mu}$; if $h$ has a chiral symmetry, then $h-\mu$ represents a class in $\operatorname{K}_{1}(A)$; if $h$ has anti-unitary symmetries, then $h-\mu$ represents a class in $\operatorname{K}_{n}(A^{\mathfrak{r}})$, the degree $n$ depends on the $\mathsf{CT}$-symmetry group of the system.
In order to compute the index pairing in a compatible way with [Bourne-Mesland:Aperiodic]*Section 6.2, we use real Kasparov modules to represent real K-theory classes, as in [Bourne-Carey-Rennie:NC_framework_topological_insulators]*Section 3.3. In particular, the real Kasparov modules in this approach consist of a finitely generated projective $A$-module, and such that the operator of this Kasparov module is zero. We remark here that the classification of $\mathsf{CT}$-symmetry groups in the setting of [Bourne-Carey-Rennie:NC_framework_topological_insulators] is different from Kellendonk, as discussed in [Kellendonk:Cst-algebraic_topological_phases]*Section 1.1.1, as it will change the degree shift in the K-theory class of Hamiltonians that have particle-hole symmetries. Since we are not concerned with the precise degree in K-theory in this article, this subtle difference will not affect our results.
@@ 2.3. Topological phases and K-theory
Let $h$ be an abstract insulator, compatible with the $\mathsf{CT}$-symmetry group $G$. Then there exist:
  • a projection $p\in\mathbb{M}_{N}(A)$;
  • a natural number $n$ and $N\in\{1,2,4\}$, together with a representation
    $\mathrm{C}\ell_{n,0}\to\mathbb{B}(pA^{\oplus N}),$
    which are determined by the $\mathsf{CT}$-symmetry group $G$ as in [Bourne-Carey-Rennie:NC_framework_topological_insulators]*table 1;
such that the following finitely generated Kasparov module represents a class $[h]\in\operatorname{KK}(\mathrm{C}\ell_{d,0},A)$:
[…]
If the system contains at most time-reversal symmetry, that is, $G\subseteq\{1,\mathsf{T}\}$, then the projection $p=p_{\mu}$ is the Fermi projection of $h$. If the system carries particle-hole symmetry or chiral symmetry, then the projection $p$ is closely related, but not equal, to $p_{\mu}$.
2.4. Position spectral triples and index pairings
In order to extact a numerical invariant from a given topological phase, described by the class $[h]\in\operatorname{KK}(\mathrm{C}\ell_{n,0},A)$ of an abstract insulator, one seeks a map
$\operatorname{KK}(\mathrm{C}\ell_{n,0},A)\to\mathbb{R}.$
One way to construct such a map is by pairing with a suitable K-homology class, represented by a (real) spectral triple:
@@ 2.4. Position spectral triples and index pairings
Let $A$ be a $\mathbb{Z}/2$-graded real C*-algebra. A (real) spectral triple $(\mathcal{A},\mathcal{H},D)$ over $A$ consists of:
@@
  • An unbounded, self-adjoint odd operator $D\colon\operatorname{dom}D\subseteq\mathcal{H}\to\mathcal{H}$
@@
  • For every $a\in\mathcal{A}$, $\varphi(a)$ maps $\operatorname{dom}D$ into $\operatorname{dom}D$; and $[D,\varphi(a)]$ extends to a bounded operator on $\mathcal{H}$.
Given $p>0$, we say a spectral triple $(\mathcal{A},\mathcal{H},D)$ is $p$-summable, if $(1+D^{2})^{-s/2}$ is trace class for all $s>p$.
A real spectral triple represents a class in real K-homology. More precisely, Kasparov has already defined his bivariant theory for “real” C*-algebras in [Kasparov:K-functor_extension]. Namely, given two $\mathbb{Z}/2$-graded “real” C*-algebras $A$ and $B$, Kasparov defined the KKR-group $\operatorname{KKR}(A,B)$, whose elements are equivalence class of “real” Kasparov $A$-$B$–modules. Restricting to the real elements $A_{\mathbb{R}}$ and $B_{\mathbb{R}}$ of $A$ and $B$ yields a real Kasparov module of $\operatorname{KKO}(A_{\mathbb{R}},B_{\mathbb{R}})$, which has the same definition as a (complex) Kasparov module, but with all Hilbert spaces and C*-algebras being real. As explained in [Bourne-Kellendonk-Rennie:Cayley_transform]*Definition 2.8, restricting to real parts does not lose any information of a “real” Kasparov module, and $\operatorname{KKR}(A,B)\simeq\operatorname{KKO}(A_{\mathbb{R}},B_{\mathbb{R}})$.
A real spectral triple over $A$ is the same thing as a real unbounded Kasparov $(A,\mathbb{R})$-module. An unbounded (real) Kasparov module is, via the bounded transform, mapped to a Fredholm module, which represents a class in K-homology. The Kasparov product between a K-theory class and a K-homology class is also referred to as an index pairing. The range is the K-theory of $\mathbb{R}$ with a degree shift, which can equal $\mathbb{Z}$, $\mathbb{Z}/2$ or 00. Thus a real spectral triple for a real C*-algebra $A$ generates a number from a topological phase in $A$ via the Kasparov product.
We note that the Kasparov product is functorial in the following sense:
@@
Let $f\colon A\to B$ be a *-homomorphism between real C*-algebras. Let $x$ be a class in $\operatorname{KK}(B\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Let $f^{*}x$ be the Kasparov product of $f\in\operatorname{KK}(A,B)$ and $x\in\operatorname{KK}(B\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Then the following diagram commutes:
${\operatorname{KK}(\mathrm{C}\ell_{d,0},A)}$${\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})}$${\operatorname{KK}(\mathrm{C}\ell_{d,0},B)}$$\scriptstyle{f^{*}x}$$\scriptstyle{f}$$\scriptstyle{x}$
@@
We call the class $f^{*}x$ the pullback of $x\in\operatorname{KK}(B\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ along $f$. When $x$ is given by a real spectral triple $(\mathcal{B},\mathcal{H},D)$, and $B$ is represented on $\mathcal{H}$ by $\varphi\colon B\to\mathbb{B}(\mathcal{H})$, then the *-algebra $\mathcal{A}=f^{-1}(\mathcal{B})$ is dense in $\mathcal{A}$, and the class of $f^{*}x$ is represented by the spectral triple $(\mathcal{A},\mathcal{H},D)$ where $A$ is represented on $\mathcal{H}$ by $\varphi\circ f\colon A\to\mathbb{B}(\mathcal{H})$.
In the following, we shall look at a class of spectral triples, which we call (tight-binding) position spectral triples as they are built from the position operators on the corresponding Hilbert space of a Delone set $\Lambda$.

### v2

@@ 2. Topological phases and K-theory
When we consider physical systems with anti-unitary symmetries, like time-reversal or particle-hole symmetries, then we must represent them as anti-unitary operators on $\mathcal{H}$ which commute or anti-commute with the Hamiltonian $H$. This turns the Hilbert space $\mathcal{H}$ into a “real” Hilbert space $(\mathcal{H},\Theta)$, and the observable C*-algebra into a “real” C*-algebra $(A,\mathfrak{r})$. The latter was introduced by Kasparov [Kasparov:K-homology]*Section 1, Definition 3.
We recall the definition of “real” and real C*-algebras and Hilbert spaces. A “real” C*-algebra is sometimes referred to as a Real C*-algebra (cf. [Bourne-Kellendonk-Rennie:Cayley_transform], note the upper case R) or a $\mathrm{C}^{*,\mathrm{r}}$-algebra (cf. [Kellendonk:Cst-algebraic_topological_phases]) in the literature.
@@
A “real” involution on a $\mathbb{Z}/2$-graded Hilbert space $\mathcal{H}$ is a conjugate-linear, grading-preserving automorphism $\Theta\colon\mathcal{H}\to\mathcal{H}$ of order two. A “real” Hilbert space is a Hilbert space $\mathcal{H}$ together with a “real” involution $\Theta$ on it.
@@
Let $\Theta$ be a “real” involution on a $\mathbb{Z}/2$-graded Hilbert space. Then $\mathbb{B}(\mathcal{H})$ becomes a $\mathbb{Z}/2$-graded “real” C*-algebra with “real” structure $T\mapsto\Theta T\Theta^{*}$. A (possibly unbounded) operator on $\mathcal{H}$ is said to be “real” if it commutes with $\Theta$.
Now we recall real C*-algebras and real Hilbert spaces.
@@
If $(A,\mathfrak{r})$ is a “real” C*-algebra, then the fixed points of its real structure
@@
2.1.1. Clifford algebras
Let $\mathbb{C}\ell_{p,q}$ be the finite-dimensional $\mathbb{Z}/2$-graded “real” C*-algebra generated by $\gamma^{1},\dots,\gamma^{p},\rho^{1},\dots,\rho^{q}$, satisfying:
@@ 2.1.1. Clifford algebras
2.1.2. The $\mathbb{Z}/2$-graded “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$
Next, we describe a representation of the graded “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ following [Ewert-Meyer:Coarse_geometry]*Section 4. The grading on $\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}$ is the tensor product of the trivial grading on $\mathrm{C}^{*}(\mathbb{Z}^{d})$ with the standard grading on $\mathbb{C}\ell_{0,d}$. The “real” structure on $\mathrm{C}^{*}(\mathbb{Z}^{d})$ is given by the pointwise complex conjugation, that is,
$\mathfrak{r}(f)(n)\mathrel{\vcentcolon=}\overline{f(n)},\quad f\in\mathrm{C}_{\mathrm{c}}(\mathbb{Z}^{d}),\;n\in\mathbb{Z}^{d}.$
@@
The Fourier transform maps the “real” C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})$ to the “real” C*-algebra $\mathrm{C}(\mathbb{T}^{d})$. Here $\mathbb{T}^{d}$ is a “real” $d$-torus, with involution $z\mapsto\overline{z}$.
@@
$\left\lvert k\right\rangle\otimes e_{i_{1}}\wedge\dots\wedge e_{i_{l}}\mapsto\frac{z^{k}}{z_{i_{1}}\dots z_{i_{l}}}\mathrm{d}z_{i_{1}}\wedge\dots\wedge\mathrm{d}z_{i_{l}},$
@@
$\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d}\simeq\mathrm{C}(\mathbb{T}^{d})\otimes\mathbb{C}\ell_{0,d}\simeq\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d}))$
where $\mathbb{C}\ell(\mathbb{T}^{d})$ is the “real” Clifford bundle of the “real” manifold $\mathbb{T}^{d}$. In particular, under this isomorphism, the canonical representation of $\mathbb{C}\ell_{0,d}$ as in ˜2.3 is mapped to the Clifford multiplication of $\mathbb{C}\ell(\mathbb{T}^{d})$ on $L^{2}\left(\bigwedge\nolimits^{*}\mathrm{T}^{*}\mathbb{T}^{d}\right)$.
We note that the “real” structure $\mathfrak{r}$ on the “real” C*-algebra $\mathrm{C}(\mathbb{T}^{d})$ induced by the involution $z\mapsto\overline{z}$ is given by
$(\mathfrak{r}f)(x)\mathrel{\vcentcolon=}\overline{f(-x)},$
thus the fixed point algebra is
$\mathrm{C}(\mathbb{T}^{d})^{\mathfrak{r}}\mathrel{\vcentcolon=}\left\{f\in\mathrm{C}(\mathbb{T}^{d})\;\middle|\;\overline{f(z)}=f(\overline{z})\right\}.$
Under the Fourier transform, $\mathrm{C}(\mathbb{T}^{d})^{\mathfrak{r}}$ is identified with the real group C*-algebra $\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$, cf. Section˜2.3.1.
Let ${\scriptstyle\mathsf{X}}_{j}$ be the $j$-th position operator on $\ell^{2}(\mathbb{Z}^{d})$ defined by
${\scriptstyle\mathsf{X}}_{j}\phi(x)\mathrel{\vcentcolon=}x_{j}\phi(x),\quad\phi\in\mathrm{C}_{\mathrm{c}}(\mathbb{Z}^{d}),\;x=(x_{1},\dots,x_{d})\in\mathbb{Z}^{d}.$
Thus the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ is an odd, essentially self-adjoint operator on the $\mathbb{Z}/2$-graded “real” Hilbert space $\ell^{2}(\mathbb{Z}^{d})\otimes\bigwedge\nolimits^{*}\mathbb{C}^{d}$.
@@
by the standard representation of $\mathbb{C}\ell_{d,0}$ in ˜2.3. The generators $\gamma^{j}$ are “real” in $\mathbb{C}\ell_{d,0}$, hence maps $\bigwedge\nolimits^{*}\mathbb{R}^{d}$ to $\bigwedge\nolimits^{*}\mathbb{R}^{d}$. The $j$-th position operator ${\scriptstyle\mathsf{X}}_{j}$ preserves $\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}$ because ${\scriptstyle\mathsf{X}}_{j}\left\lvert k\right\rangle=k_{j}\left\lvert k\right\rangle$ and $k_{j}$ is real for any $k\in\mathbb{Z}^{d}$. Therefore the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ preserves $\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$. ∎
2.2. Kasparov theory for real C*-algebras
In his seminal work [Kasparov:K-functor_extension], Kasparov introduced a bivariant K-theory for $\mathbb{Z}/2$-graded C*-algebras. Elements in his bivariant K-theory group $\operatorname{KK}(A,B)$ may be viewed as “generalised homomorphisms” from $A$ to $B$. Kasparov’s constructions are general enough to be transferred to the setting of real C*-algebras (cf. [Bourne-Carey-Rennie:NC_framework_topological_insulators]*Appendix A).
@@ 2.2. Kasparov theory for real C*-algebras
(2.7) $\operatorname{KO}_{n}(B)\simeq\operatorname{KK}(\mathrm{C}\ell_{n,0},B).$
Given $\mathbb{Z}/2$-graded real C*-algebras $A_{1},A_{2},B_{1},B_{2},D$, there is a natural group homomorphism
$\times_{D}\colon\operatorname{KK}(A_{1},B_{1}\otimes D)\times\operatorname{KK}(A_{2}\otimes D,B_{2})\to\operatorname{KK}(A_{1}\otimes A_{2},B_{1}\otimes B_{2})$
called the Kasparov product, cf. [Kasparov:K-functor_extension]*Section 4, Theorem 4. We shall need the following formal constructions with the Kasparov product, and refer to [Kasparov:K-functor_extension, Bourne-Carey-Rennie:NC_framework_topological_insulators] for details.
Functoriality
Let $f\colon A\to B$ and $g\colon B\to C$ be *-homomorphisms between real C*-algebras. Then $f$ gives an element $f]\in\operatorname{KK}(A,B)$ and $g]\in\operatorname{KK}(B,C)$. Moreover, we have
$f]\times_{B}[g]=[g\circ f]\in\operatorname{KK}(A,C).$
Graded $\operatorname{KO}_{*}(\mathbb{R})$-module structure of $\operatorname{KO}_{*}(B)$
Let $A_{1}=\mathrm{C}\ell_{n,0}$, $A_{2}=\mathrm{C}\ell_{m,0}$, $B_{1}=B_{2}=D=\mathbb{R}$. Then the Kasparov product gives a group homomorphism
$\times_{\mathbb{R}}\colon\underbrace{\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathbb{R})}_{\simeq\operatorname{KO}_{n}(\mathbb{R})}\times\underbrace{\operatorname{KK}(\mathrm{C}\ell_{m,0},\mathbb{R})}_{\simeq\operatorname{KO}_{m}(\mathbb{R})}\to\underbrace{\operatorname{KK}(\mathrm{C}\ell_{m+n,0},\mathbb{R})}_{\simeq\operatorname{KO}_{m+n}(\mathbb{R})},$
where the isomorphisms are given as in (2.7). This turns $\operatorname{KO}_{*}(\mathbb{R})\mathrel{\vcentcolon=}\bigoplus_{n}\operatorname{KO}_{n}(\mathbb{R})$ into a graded commutative ring.
Now let $A_{1}=\mathrm{C}\ell_{n,0}$, $A_{2}=\mathrm{C}\ell_{m,0}$, $B_{1}=D=\mathbb{R}$ and $B_{2}=B$ be an ungraded real C*-algebra. Then the Kasparov product gives a group homomorphism
$\times_{\mathbb{R}}\colon\underbrace{\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathbb{R})}_{\simeq\operatorname{KO}_{n}(\mathbb{R})}\times\underbrace{\operatorname{KK}(\mathrm{C}\ell_{m,0},B)}_{\simeq\operatorname{KO}_{m}(B)}\to\underbrace{\operatorname{KK}(\mathrm{C}\ell_{m+n,0},B)}_{\simeq\operatorname{KO}_{m+n}(B)},$
which turns $\operatorname{KO}_{*}(B)\mathrel{\vcentcolon=}\bigoplus_{n}\operatorname{KO}_{n}(B)$ into a graded commutative module over $\operatorname{KO}_{*}(\mathbb{R})$.
Index pairing
Let $A_{1}=\mathrm{C}\ell_{n,0}$, $A_{2}=\mathrm{C}\ell_{0,d}$, $B_{1}=B_{2}=\mathbb{R}$, then the Kasparov product gives a group homomorphism
$\times_{D}\colon\underbrace{\operatorname{KK}(\mathrm{C}\ell_{n,0},D)}_{\simeq\operatorname{KO}_{n}(D)}\times\operatorname{KK}(D\otimes\mathrm{C}\ell_{0,d},\mathbb{R})\to\underbrace{\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})}_{\simeq\operatorname{KO}_{n-d}(\mathbb{R})}.$
Thus every element $\alpha\in\operatorname{KK}(D\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ induces a group homomorphism
(2.8) $\times_{D}\alpha\colon\operatorname{KO}_{n}(D)\to\operatorname{KO}_{n-d}(\mathbb{R}).$
Moreover, since the Kasparov product is functorial, $\alpha$ actually induces an $\operatorname{KO}_{*}(\mathbb{R})$-module homomorphism $\operatorname{KO}_{*}(D)\to\operatorname{KO}_{*-d}(\mathbb{R})$.
The group homomorphism (2.8) induced by an element $\alpha\in\operatorname{KK}(D\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ is called an index pairing. Such an element $\alpha$ can be represented by a real spectral triple.
@@
Let $A$ be a $\mathbb{Z}/2$-graded real C*-algebra. A real spectral triple $(\mathcal{A},\mathcal{H},D)$ over $A$ consists of:
@@
  • An unbounded, self-adjoint odd operator $D\colon\operatorname{dom}D\subseteq\mathcal{H}\to\mathcal{H}$;
@@
  • For every $a\in\mathcal{A}$, $\varphi(a)$ maps $\operatorname{dom}D$ into $\operatorname{dom}D$; and the graded commutator $D,\varphi(a)]$ extends to a bounded operator on $\mathcal{H}$.
Let $\xi\mathrel{\vcentcolon=}(\mathcal{B},\mathcal{H},D)$ be a real spectral triple over a real C*-algebra $B$, where $B$ is represented on $\mathcal{H}$ via $\varphi\colon B\to\mathbb{B}(\mathcal{H})$. Then $\xi$ represents a class $\xi]\in\operatorname{KK}(B,\mathbb{R})$.
Let $f\colon A\to B$ be a *-homomorphism. Then the *-algebra $\mathcal{A}\mathrel{\vcentcolon=}f^{-1}(\mathcal{B})$ is dense in $A$, and $f^{*}\xi\mathrel{\vcentcolon=}(\mathcal{A},\mathcal{H},D)$ is a real spectral triple over $A$, where $A$ is represented on $\mathcal{H}$ via $\varphi\circ f\colon A\to\mathbb{B}(\mathcal{H})$.
We call $f^{*}\xi\mathrel{\vcentcolon=}(\mathcal{A},\mathcal{H},D)$ in the above example, the pullback spectral triple of $(\mathcal{B},\mathcal{H},D)$ along $f$. Functoriality of Kasparov product states the following:
@@
Let $f\colon A\to B$ be a *-homomorphism between real C*-algebras. Let $\xi$ be a real spectral triple which represents a class $\xi]\in\operatorname{KK}(B\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Then the pullback spectral triple $f^{*}\xi$ represents the Kasparov product $f]\times_{B}[\xi]\in\operatorname{KK}(A\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. Moreover, the following diagram commutes:
$\operatorname{KK}(\mathrm{C}\ell_{n,0},A)$$\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})$$\operatorname{KK}(\mathrm{C}\ell_{n,0},B)$$\scriptstyle{\times_{A}[f^{*}\xi]}$$\scriptstyle{\times_{A}[f]}$$\scriptstyle{\times_{B}[\xi]}$
@@
2.3. Position spectral triples
In the following, we focus on a class of spectral triples, which we call position spectral triples as they are built from the position operators on the corresponding Hilbert space of a discrete set $\Lambda$.